# Collision-Free Navigation of Wheeled Mobile Robots: An Integrated Path Planning and Tube-Following Control Approach

### v1

@@ IV Simulation Results
Figure 6: Navigation trajectories of the two path planners.
Figure 7: Velocity norms of the two path planners.
We next illustrate the performance of the proposed adaptive tube-following controller (29) (denoted as “TF controller”) in the presence of disturbance given by (47). As a case study, the trajectory starting from the 3nd initial position is selected as the reference trajectory (see Fig. 6). For comparison purposes, the extensively used proportional-integral (PI) controller of the following form is simulated:
$\mathbf{u}=-k_{p}\mathbf{x}_{e}-k_{i}\int_{0}^{t}\mathbf{x}_{e}(\tau)\text{d}\tau,$
where $k_{p}=0.5$ and $k_{i}=0.3$ are constant gains. Furthermore, in order to verify the robustness of the PF-based planner (48) (denoted here as “PF controller”), we execute it in the presence of disturbances and define the difference between the ideal and actual trajectories as the position tracking error $\mathbf{x}_{e}$. The closed-loop responses of these three controllers are shown in Fig. 8. In the top left subfigure of Fig. 8, the tracking error $\mathbf{x}_{e}$ under both the TF and PI controllers is shown to remain within a small neighborhood of the origin (but quantitatively, the TF controller has higher tracking accuracy than the PI controller), while satisfying the tube constraint (24). However, the PF controller yields a larger tracking error and violates the tube constraint. The latter may result in collision with obstacles or the workspace boundary. The WMR’s heading angle is plotted in the top right subfigure of Fig. 8, from which it is clear that $\theta$ under the TF and PI controllers is bounded and converges to a nearly constant value (precisely, the TF controller has a higher stability), while under the PF controller $\theta$ continuously increases. This implies that the PF controller (i.e., path planner (48)) is susceptible to disturbances. The control input norm $\|\mathbf{u}\|$ is shown in the bottom left subfigure of Fig. 8. Obviously, the control input $\mathbf{u}$ of the TF and PI controllers satisfies the magnitude limits $\|\mathbf{u}\|\leq u_{m}$, while the PF controller fails. The parameter estimate $\hat{d}$ is given in the bottom right subfigure of Fig. 8. As can be seen, $\hat{d}$ remains within the set $\Omega_{\delta}$, indicating the effectiveness of the projection adaptive law (31). Finally, the position tracking trajectory of the proposed TF controller (29) is shown in Fig. 9. Intuitively, the virtual control point $\bm{P}$ accurately tracks the reference trajectory, while remaining with the predefined tube. Therefore, the WMR reaches the goal position without colliding with any obstacles (dark gray balls), as clearly seen in Fig. 9. Overall, the proposed integrated path planning and tube-following control scheme can successfully drive the WMR to reach the predefined goal without colliding with any obstacles, despite the presence of input disturbances.
Figure 8: Closed-loop control responses.
@@ V Experiments
In this section, several experiments are carried out to verify the effectiveness of the proposed method.
V-A Experimental setup
@@
In the experiments, several simple scenarios are considered where the WMR is tasked with navigating towards a designated goal location, while ensuring avoidance of any obstacles en route. The experimental platform is shown in Fig. 10, which consists of a Mona robot [31] with nonholonomic dynamics, several hot obstacles containing iron powders that can heat up themselves, a RGB camera to observe the pose and location of the WMR, a thermal camera to obtain the size and location of the hot obstacles, and a control PC to collect feedback data and send motion commands. The Wi-Fi communication (at a rate of $10\,\rm Hz$) between the control PC and the robot is built by ROS. All experiments are performed in an arena with a size of $\rm 1.8\,m\times 0.8\,m$. It should be emphasized that since the WMR in our lab has a very limited sensing capability, the distance between the WMR and obstacle boundary is obtained in real-time from the RGB and thermal cameras, rather than sensors onboard the WMR. Regarding the robot configuration, we set the radius of the circle surrounding the robot to $r=0.06\,\text{m}$, the parameter for the change of coordinates to $\ell=-0.02$, the safety margin to $\epsilon=0.04\,\text{m}$, and the obstacle influence region to $\epsilon^{*}=0.05\,\text{m}$. In all experiments, the parameters of the path planner are set to $\alpha=0.03$ and $\beta=0.005$, while the controller parameters are set to $\rho=0.03$, $k=1.5$, $\psi=0.005$, $\eta=0.1$, $\gamma=0.01$, $d_{m}=0.036$, $\delta=0.005$, and $d(0)=0.01$.
V-B Experimental results
We here report the results of two representative experiments, where the extracted thermal image and the RGB image are merged to record the experimental process. In the first experiment, six obstacles are placed in the arena. The WMR starts from the left side of the arena (coordinates: $[0.27,0.21]^{\top}\text{m}$), and is tasked with reaching the goal location situated on the right side (coordinates: $[1.73,0.54]^{\top}\text{m}$). The results are shown in Fig. 11, from which it can be observed that the proposed path planner generates a collision-free navigation path for the WMR (i.e., the red line in Fig. 11(a)), and the tube-following controller guarantees that the WMR tracks the desired path along with the predefined safe tube. The latter can be seen from Fig. 11(c), where the norm of the position tracking error is smaller than the tube width $\rho=0.03\text{m}$ all the time. In the second experiment, we add two more hot obstacles to the arena, which makes the environment more complex for the WMR. In this case, the WMR starts from the right side of the arena (coordinate: $[1.83,0.68]^{\top}\text{m}$) and is required to reach a goal location at the left side (coordinate: $[0.20,0.40]^{\top}\text{m}$). The experiment results are shown in Fig. 12, where we can observe similar outcomes as in the first experiment. Our path planner successfully generates a collision-free navigation path for the WMR and the tube-following controller drives the WMR to track the generated path without running out the predefined tube (see Fig. 12(c)). Furthermore, we perform multiple trials with different initial and goal locations to further test the proposed method. From these experiments, we consistently observe the effectiveness of our proposed robot navigation algorithm. More details can be found in the supplementary video (https://vimeo.com/895801720).
Figure 10: The experiment setup: (a) experimental platform; (b) cameras; (c) Mona robot.
Figure 11: The experimental scenario with six obstacles: (a) the generated path (red line) and the actual trajectory (black line) of the WMR; (b) screenshots of the experiment; (c) position error; (d) WMR’s orientation; (e) parameter estimation $\hat{d}$.
Figure 12: The experimental scenario with eight obstacles: (a) the generated path (red line) and the actual trajectory (black line) of the WMR; (b) screenshots of the experiment; (c) position error; (d) WMR’s orientation; (e) parameter estimation $\hat{d}$.
@@ VI Conclusion
In this paper, we have proposed an integrated path planning and tube-following control scheme for collision-free navigation of WMRs in a convex workspace cluttered with multiple spherical obstacles. A path planner was designed using Bouligand’s tangent cones to generate safe, collision-free motion paths for the WMR. The resulting vector field asymptotically guides the WMR from almost all initial configurations in the free space towards the goal location, without colliding with any obstacles along the way. Then, an adaptive tube-following controller was derived, which guarantees that the actual trajectory of the WMR remains within a predefined safe tube around the reference trajectory, despite the presence of disturbances. The tube-following control significantly enhances the safety of robot navigation in the perturbed scenarios. In particular, the proposed closed-loop navigation strategy can be implemented using only the local sensing information. Finally, numerical simulations and experiments illustrate the effectiveness of the proposed autonomous navigation algorithm. Future work will focus on reactive collision-free navigation of multiple WMRs. [Proof of Theorem 2] From (18) and (19), it follows that $\bm{\tau}(\mathbf{x})\in\mathbf{T}_{\mathcal{X}_{\epsilon}}(\mathbf{x})$ for all $\mathbf{x}\in\mathcal{X}_{\epsilon}$, which according to Theorem 1 suggests that the free space $\mathcal{X}_{\epsilon}$ is forward invariant. As previously stated, $\bm{\tau}(\mathbf{x})$ is piecewise continuously differentiable, indicating that it is locally Lipschitz on its domain $\mathcal{X}$ [32]. Moreover, the free space $\mathcal{X}{\epsilon}$, as a compact subset of $\mathcal{X}$, is forward invariant. Then, it follows from [33, Theorem 3.3] that the closed-loop kinematics (11) with $\mathbf{x}(0)\in\mathcal{X}_{\epsilon}$ has a unique solution that is defined for all $t\geq 0$. Subsequently, we show that the unique solution converges to a set of stationary points. To this end, we consider the following LFC:
@@
$\bm{\tau}(\mathbf{x})=\underbrace{(c(\mathbf{x})\mathbf{I}_{2}+(1-c(\mathbf{x}))\mathbf{\Pi}(\mathbf{x}))}_{\mathbf{\Theta}(\mathbf{x})}\bm{\kappa}_{0}(\mathbf{x}),$ (50)
@@
$\dot{W}(\mathbf{x})=-k_{0}(\mathbf{x})(\mathbf{x}-\mathbf{x}^{\ast})^{\top}\mathbf{\Theta}(\mathbf{x})(\mathbf{x}-\mathbf{x}^{\ast}).$ (52)
Since $k_{0}(\mathbf{x})>0$ and $\mathbf{\Pi}(\mathbf{x})$ is positive semidefinite, one has
@@
indicating that $\mathbf{x}=\mathbf{x}^{\ast}$ is a stable equilibrium of (11). From LaSalle’s invariance principle, it follows that the solution of (11) asymptotically converges to the set $\{\mathbf{x}\in\mathcal{X}_{\epsilon}\mid\dot{W}(\mathbf{x})=0\}$, i.e., a set of stationary points. One can infer from (19) and (52) that the stationary points correspond to either $\mathbf{x}=\mathbf{x}^{\ast}$ or the points satisfying $\text{d}_{\mathcal{O}}(\mathbf{x})=\epsilon$ and $\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{b}(\mathbf{x})/\|\bm{\kappa}_{0}(\mathbf{x})\|=1$. Therefore, for any $\mathbf{x}(0)\in\mathcal{X}_{\epsilon}$, the solution of (11) converges to the set $\{\mathbf{x}^{\ast}\}\cup_{i=1}^{n}\mathcal{S}_{i}$. The set contains $n+1$ stationary points, all of which are isolated due to Assumption 2. Geometrically, the undesired stationary point $\mathbf{s}_{i}\in\mathcal{S}_{i}$ (on the boundary of $\mathcal{O}_{i}^{\epsilon}$) and $\mathbf{x}^{\ast}$ are collinear with $\mathbf{c}_{i}$, but located on the opposite sides of $\mathbf{c}_{i}$, as seen in Fig. 13. As per (52), the analytical form of $\mathbf{s}_{i}$ is obtained as follows:
@@
The isolated stationary points $\mathbf{s}_{i}$, $i\in\mathbb{I}$ may prevent us from achieving the objective (motion-to-goal). To check the stability of $\mathbf{s}_{i}$, we define a small neighborhood of $\mathbf{s}_{i}$, i.e., an open ball $\mathcal{B}(\mathbf{s}_{i},\epsilon^{\prime})$ with $0<\epsilon^{\prime}\leq\min\{\epsilon,\epsilon^{\ast}-\epsilon\}$, such that $\bf{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{b}(\mathbf{x})>0$ holds for all $\mathbf{x}\in\mathcal{B}(\mathbf{s}_{i},\epsilon^{\prime})$. When restricted on $\mathcal{B}(\mathbf{s}_{i},\epsilon^{\prime})$, the feasible set of initial configurations is $\mathcal{F}_{i}:=\mathcal{X}_{\epsilon}\cap\mathcal{B}(\mathbf{s}_{i},\epsilon^{\prime})$, and $\mathbf{x}$ evolves according to
$\dot{\mathbf{x}}=-k_{0}(\mathbf{x})(\mathbf{I}_{2}-\phi(\text{d}_{\mathcal{O}}(\mathbf{x}))\mathbf{b}(\mathbf{x})\mathbf{b}^{\top}(\mathbf{x}))(\mathbf{x}-\mathbf{x}^{\ast}).$ (55)
@@
Geometrically speaking, all elements of $\mathcal{M}_{i}$ form a radial line extending outward from the stationary point $\mathbf{s}_{i}$ along $\mathbf{s}_{i}-\mathbf{c}_{i}$ (see the yellow line in Fig. 13), while all elements of $\mathcal{N}_{i}$ form a curve, which is a portion of the boundary $\partial\mathcal{O}_{i}^{\epsilon}$ (see the blue curve in Fig. 13). In the following, three cases are considered: 1) $\mathbf{x}(0)\in\mathcal{B}(\mathbf{s}_{i},\epsilon^{\prime})\cap\mathcal{M}_{i}$; 2) $\mathbf{x}(0)\in\mathcal{N}_{i}\setminus\{\mathbf{s}_{i}\}$; and 3) $\mathbf{x}(0)\in\mathcal{F}_{i}\setminus(\mathcal{M}_{i}\cup\mathcal{N}_{i})$. Case 1): $\mathbf{x}(0)\in\mathcal{B}(\mathbf{s}_{i},\epsilon^{\prime})\cap\mathcal{M}_{i}$. Note that for all $\mathbf{x}\in\mathcal{M}_{i}$, $\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{b}(\mathbf{x})/\|\bm{\kappa}_{0}(\mathbf{x})\|=1$ holds, whereby (55) reduces to
$\dot{\mathbf{x}}=-k_{0}(\mathbf{x})(1-\phi(\text{d}_{\mathcal{O}}(\mathbf{x})))(\mathbf{x}-\mathbf{x}^{\ast}).$ (58)
From (20) and (56), it is evident that $0\leq\phi(\text{d}_{\mathcal{O}}(\mathbf{x}))\leq 1$ on the set $\mathcal{M}_{i}$ and $\phi(\text{d}_{\mathcal{O}}(\mathbf{x}))=1$ only when $\mathbf{x}=\mathbf{s}_{i}$. Moreover, $k_{0}(\mathbf{x})$ is strictly larger than 00 by construction. Hence, for all $\mathbf{x}\in\mathcal{M}_{i}\setminus\{\mathbf{s}_{i}\}$, the control vector $\bm{\tau}(\mathbf{x})$ (i.e., the right-hand side of (58)) points directly towards $\mathbf{s}_{i}$ and becomes zero when $\mathbf{x}=\mathbf{s}_{i}$, indicating that the set $\mathcal{M}_{i}$ is forward invariant. Consider the following LFC:

### v2

@@ IV Simulation Results
We begin by illustrating the effectiveness of the proposed prescribed-time planner (19) (denoted as “PTP”). For comparison purposes, the APF-based planner presented in [32] (denoted as “APF”) and the CBF-based planner presented in [9] (denoted as “CBF”) are also simulated. The attractive and repulsive potentials of the APF planner are designed as $U_{att}(\mathbf{x}_{d})=\frac{k_{a}}{2}\|\mathbf{x}_{d}-\mathbf{x}^{\ast}\|^{2}$, $U_{rep}(\mathbf{x}_{d})=k_{r}\sum_{i=1}^{8}\Upsilon(\text{d}_{\mathcal{O}_{i}}(\mathbf{x}_{d}))$, where $k_{a}$ and $k_{r}$ are positive weights, and $\Upsilon$ is a smooth repulsive function defined as $\Upsilon(z)=(\epsilon^{\ast}-z)^{2}\text{ln}(z-\epsilon)/(z-\epsilon)$ if $z\in(\epsilon,\epsilon^{\ast}]$, and $\Upsilon(z)=0$, if $z>\epsilon^{\ast}$. Then, the vector field generated by the APF planner is
We consider a CBF $f(\mathbf{x}_{d}):=\min_{i\in\{0\}\cup\mathbb{I}}f_{i}(\mathbf{x}_{d})$ together with the safe set :=$\mathcal{F}:=\{\mathbf{x}_{d}\in\mathbb{R}^{2}\mid f(\mathbf{x}_{d})\geq 0\}$, where $f_{0}(\mathbf{x}_{d}):=1-(x_{d}/2.9)^{20}-(y_{d}/1.4)^{20}$ (for workspace boundaries), and $f_{i}(\mathbf{x}_{d}):=\|\mathbf{x}_{d}-\mathbf{c}_{i}\|^{2}-(r+r_{i}+\epsilon)^{2}$, $i\in\mathbb{I}$. The CBF planner is obtained by quadratic programming
$\mathop{\rm argmin}\limits_{\bm{\tau}_{d}\in\mathbb{R}^{2}}$ $~{}\|\bm{\tau}_{d}-\bm{\tau}_{d,\rm des}(\mathbf{x}_{d})\|^{2}$ (CBF-QP)
s.t. $~{}\nabla f(\mathbf{x}_{d})^{\top}\bm{\tau}_{d}\geq-\gamma f(\mathbf{x}_{d}),$
where $\alpha>0$ is a user-defined constant. The CBF-QP has an explicit solution of the form
$\bm{\tau}_{d}=\bm{\tau}_{d,\rm des}(\mathbf{x}_{d})+\bm{\tau}_{d,\rm safe}(\mathbf{x}_{d}),$ (40)
where $\bm{\tau}_{d,\rm des}(\mathbf{x}_{d}):=\bm{\kappa}_{0}(\mathbf{x}_{d})$, and $\bm{\tau}_{d,\rm safe}$ is determined by
$\bm{\tau}_{d,\rm safe}(\mathbf{x}_{d}):=\left\{\begin{aligned} &-\dfrac{\nabla f(\mathbf{x}_{d})}{\nabla f(\mathbf{x}_{d})^{\top}\nabla f(\mathbf{x}_{d})}\Psi(\mathbf{x}_{d}),&~{}&\text{if}~{}\Psi<0,\\ &0,&~{}&\text{if}~{}\Psi\geq 0,\end{aligned}\right.$
with $\Psi(\mathbf{x}_{d}):=\nabla f(\mathbf{x}_{d})^{\top}\bm{\tau}_{d,\rm des}(\mathbf{x}_{d})+\gamma f(\mathbf{x}_{d})$. As dictated by [9], the CBF planner (40) ensures that the set $\mathcal{F}$ is forward invariant. The APF planner parameters are chosen as $k_{a}=k_{0}$ and $k_{r}=0.1$, while the CBF planner parameter is chosen as $\gamma=0.1$. It is important to note that these three planners have the same motion-to-goal control law. Their design parameters are listed in Table II.
TABLE II: Planner and controller parameters.
Method Parameter
PTP in (19) $k_{0}=0.01$, $T=200$, $\varsigma=0.5$
TFC in (26) $\rho=0.06$, $k_{1}=0.8$, $k_{2}=0.001$, $T_{f}=200$, $\varsigma_{f}=3$
APF in [32] $k_{a}=k_{0}=0.01$, $k_{r}=0.1$
CBF in [9] $\gamma=0.1$
The planned trajectories starting at a set of initial positions (purple points) are plotted in Fig. 6, where the dense arrows denote the magnitude and direction of the resulting vector field from the prescribed-time planner (19) at various points. These vector arrows are directly obtained as per (19). As shown in Fig. 6, all three planners successfully guide the WMR from different initial positions to the goal (red point), while avoiding all obstacle regions. The comparison results of $\|\mathbf{x}_{d}\|$ and $\|\bm{\tau}_{d}\|$ are depicted in Fig. 7, from which we find that both the APF and CBF converge slower than the proposed PTP; moreover, their convergence times become longer as the path is longer. However, the PTP always converges at $T=200\,\rm s$, regardless of initial positions. We further provide quantitative comparison results in Table III, where the 3rd initial position is considered as a case study. Although all the three planners offer analytical solutions and have an almost identical time consumption (average time of 100 runs), the convergence time differs greatly. The PTP converges at the prescribed time $T=200\,\rm s$, while the APF and CBF take significantly longer, converging in $950\,\rm s$. The path length metric also differs, with CBF covering a shorter path compared to the other two methods. The maximum and standard deviation (std) of the velocity norm provide additional insights, with PTP exhibiting higher values than APF and CBF in both measures, due to its faster convergence rate.
Figure 6: Planned trajectories of the three path planners.
(a) PTP in our work
(b) APF in [32]
(c) CBF in [9]
Figure 7: Control responses of different path planners.
TABLE III: Planning performance comparison.
Method Time consumption Convergence time Path length Maximum of $\|\bm{\tau}_{d}\|$ std of $\|\bm{\tau}_{d}\|$ Obstacle avoidance Prescribed time
PTP in (19) 0.3136 s 200 s 6.4411 m 0.0548 0.0148
APF in [32] 0.3196 s 950 s 6.7394 m 0.0548 0.0096
CBF in [9] 0.3474 s 950 s 6.3976 m 0.0479 0.0102
In the following, we show the effectiveness and performance of the proposed tube-following controller (26) (denoted as “TFC”) with parameters given in Table II. As a case study, the planned trajectory starting from the 3rd initial position is selected as the reference trajectory. To verify the robustness of the APF and CBF planners, we execute them in the presence of disturbances and define the difference between the reference and actual trajectories as the position tracking error $\mathbf{x}_{e}$. The closed-loop tracking responses of three controllers are shown in Fig. 8. In Fig. 8(a), the tracking error $\mathbf{x}_{e}$ under the proposed TFC is shown to converge to a residual set of $3.74\times 10^{-4}\,\rm m$ after the prescribed time $T_{f}=200\,\rm s$, while complying with the tube constraint defined by (22). In contrast, the APF and CBF controllers exhibit larger stead-state errors of $0.43\,\rm m$ and $0.32\,\rm m$, respectively, whilst violate the tube constraint, leading to an increased risk of collision. The WMR’s heading angle is provided in Fig. 8(b), from which it is clear that $\theta$ under the TFC is bounded and converges to a nearly constant value, while under the APF and CBF, $\theta$ continuously increases. This implies that both the APF and CBF controllers are susceptible to disturbances. The control norm $\|\mathbf{u}\|$ is shown in Fig. 8(c). To further show the prescribed-time convergence property of the proposed TFC, the tracking error norm under different prescribed time $T_{f}$ is given in Fig. 8(d). As illustrated, the proposed controller can precisely and flexibly set the convergence time in advance. The position tracking trajectory of the proposed TFC is shown in Fig. 9. Intuitively, $\mathbf{x}$ accurately tracks the reference trajectory, while remaining within the safe tube. This implies that, under the proposed InPTC, the WMR reaches the goal within the prescribed time $T=200\,\rm s$, while avoiding collision with any obstacles (dark gray balls).
Figure 8: Comparisons of the closed-loop tracking responses.
@@ V Experiments
The experimental platform is shown in Fig. 10, which consists of a Mona robot [33] with nonholonomic dynamics, several hot obstacles containing iron powders that can heat up themselves, a RGB camera to observe the pose and location of the WMR, a thermal camera to obtain the size and location of the hot obstacles, and a control PC to collect feedback data and send motion commands. The Wi-Fi communication (at a rate of $10\,\rm Hz$) between the control PC and robot is built by ROS. All experiments are performed in a $\rm 2.78\,m\times 1.4\,m$ arena. Since the Mona robot has a very limited onboard sensing capability, the distance between the robot and obstacle boundary is obtained in real-time from the RGB and thermal cameras. The radius of the circle surrounding the robot is set to $r=0.06\,\text{m}$, the parameter for the change of coordinates is $\ell=-0.02\,\rm m$. The design parameters of the InPTC are listed in Table IV. In experiments, the disturbances could come from measurement noises, communication delay, imperfect wheel alignment, motor or actuator asymmetry, etc. In the following, we report the results of two representative experiments, where the extracted thermal image and the RGB image are merged to record the experimental process.
Figure 10: The experiment setup: (a) platform; (b) cameras; (c) Mona robot.
TABLE IV: Design parameters of the InPTC.
Method Parameter
Planner in (19) $k_{0}=0.01$, $T=250$, $\varsigma=0.5$
Controller in (26) $\rho=0.06$, $k_{1}=0.01$, $k_{2}=0.003$, $T_{f}=150$, $\varsigma_{f}=2$
In the first experiment, eight circular obstacles are placed in the arena, and the safety margin and influence region of obstacles are set to $\epsilon=0.04\,\text{m}$ and $\epsilon^{*}=0.05\,\text{m}$, respectively. The WMR starts from the right side of the arena (coordinates: $[2.45,1.27]^{\top}\text{m}$), and is tasked with reaching the goal located on the right side (coordinates: $[0.44,0.35]^{\top}\text{m}$). The task is required to be completed within $250\,\rm s$. The results are shown in Fig. 11, from which it is clear that the proposed planner in (19) generates a smooth, collision-free reference trajectory that converges to the goal at the preassigned time $T=250\,\rm s$ (i.e., the red line in Fig. 11(a)). While the proposed controller in (26) achieves trajectory tracking along with the safe tube (see Fig. 11(c)), with a tracking accuracy higher than $4.9\times 10^{-3}\,\rm m$ after the preassigned time $T_{f}=150\,\rm s$ (see Fig. 11(d)). As seen in Fig. 11(e), the robot’s heading angle remains nearly constant after the task completion time $T=250\,\rm s$. The above results demonstrate that the proposed InPTC scheme enables the WMR to safely accomplish the navigation task within the required time of $250\,\rm s$.
In the second experiment, we consider a more complex environment with six obstacles of various shapes. The safety margin and influence region of obstacles are set to $\epsilon=0.08\,\text{m}$ and $\epsilon^{*}=0.1\,\text{m}$, respectively. The WMR starts from the left side of the arena (coordinate: $[0.3,0.5]^{\top}\text{m}$) and is required to reach a goal location at the right side (coordinate: $[2.5,1.0]^{\top}\text{m}$). The results are depicted in Fig. 12, hich shows outcomes analogous to those observed in the initial experiment. In this case, the proposed InPTC still generates a smooth and safe reference trajectory with prescribed-time convergence (see Fig. 12(c)). It also achieves prescribed-time tube-following control with an accuracy of $4.7\times 10^{-3}\,\rm m$ (see Fig. 12(d)), thus enabling the WMR to accomplish the navigation task within the prescribed time $T=250\,\rm s$ (see Fig. 12(a)).
Furthermore, we perform multiple trials with different initial and goal locations to further test the proposed InPTC. From these experiments, we consistently observe the effectiveness of InPTC in achieving tracking accuracy higher than $0.01\,\rm m$ and task completion time less than $250\,\rm s$. More details are provided in the supplementary video (https://vimeo.com/895801720).
Figure 11: The experiment with circular obstacles: (a) WMR’s motion trajectory; (b) screenshots of the experiment; (c) reference position norm; (d) position error norm; (e) WMR’s heading angle.
Figure 12: The experiment with circular and polygonal obstacles: (a) WMR’s motion trajectory; (b) screenshots of the experiment; (c) reference position norm; (d) position error norm; (e) WMR’s heading angle.
@@ VI Conclusion
In this article, we propose a prescribed-time path planning algorithm for wheeled mobile robots (WMRs) to generate a collision-free reference trajectory that can converge to the goal location at a prescribed finite time. The derived planner is then paired with a tube-following controller that allows WMRs to achieve high-precision trajectory tracking within a prescribed time, while remaining within a predefined “safe tube” around the reference trajectory. The resulting integrated planning and tube-following control (InPTC) scheme achieves prescribed-time collision-free navigation of WMRs, despite the presence of disturbances. The salient features of the proposed InPTC are two-fold: 1) it enables WMRs to accomplish safe navigation within a finite time that can be flexibly preassigned according to the mission requirements; and 2) it enhances the navigation safety of WMRs in practical implementations. Simulation and experimental results show the effectiveness of InPTC. Future work will focus on extending the proposed method to achieve prescribed-time safe navigation of WMRs in environments with dynamic or intersecting obstacles.
[Proof of Theorem 2]
From (11) and (12), it follows that $\mathbf{h}(\mathbf{x})\in\mathbf{T}_{\mathcal{X}_{\epsilon}}(\mathbf{x})$ for all $\mathbf{x}\in\mathcal{X}_{\epsilon}$, which according to Theorem 1 suggests that the free space $\mathcal{X}_{\epsilon}$ is forward invariant.
As previously stated, $\mathbf{h}(\mathbf{x})$ is piecewise continuously differentiable, indicating that it is locally Lipschitz on its domain $\mathcal{X}$ [34]. Moreover, the free space $\mathcal{X}{\epsilon}$, as a compact subset of $\mathcal{X}$, is forward invariant. Then, it follows from [35, Theorem 3.3] that the closed-loop kinematics (7) under $\bm{\tau}=\mathbf{h}(\mathbf{x})$ with $\mathbf{x}(0)\in\mathcal{X}_{\epsilon}$ has a unique solution that is defined for all $t\geq 0$. We next show that the unique solution converges to a set of stationary points. To this end, we consider the following LFC:
@@
$\mathbf{h}(\mathbf{x})=\underbrace{(c(\mathbf{x})\mathbf{I}_{2}+(1-c(\mathbf{x}))\mathbf{\Pi}(\mathbf{x}))}_{\mathbf{\Theta}(\mathbf{x})}\bm{\kappa}_{0}(\mathbf{x}),$ (42)
@@
$\dot{W}(\mathbf{x})=-k_{0}(\mathbf{x}-\mathbf{x}^{\ast})^{\top}\mathbf{\Theta}(\mathbf{x})(\mathbf{x}-\mathbf{x}^{\ast}).$ (44)
Since $k_{0}>0$ and $\mathbf{\Pi}(\mathbf{x})$ is positive semidefinite, one has
@@
indicating that $\mathbf{x}=\mathbf{x}^{\ast}$ is a stable equilibrium of (7). From LaSalle’s invariance principle, it follows that the solution of (7) asymptotically converges to the set $\{\mathbf{x}\in\mathcal{X}_{\epsilon}\mid\dot{W}(\mathbf{x})=0\}$, i.e., a set of stationary points. One can infer from (12) and (44) that the stationary points correspond to either $\mathbf{x}=\mathbf{x}^{\ast}$ or the points satisfying $\text{d}_{\mathcal{O}}(\mathbf{x})=\epsilon\wedge\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{b}(\mathbf{x})/\|\bm{\kappa}_{0}(\mathbf{x})\|=1$. Therefore, for any $\mathbf{x}(0)\in\mathcal{X}_{\epsilon}$, the solution of (7) converges to the set $\mathcal{C}=\{\mathbf{x}^{\ast}\}\cup_{i=1}^{n}\{\mathbf{s}_{i}\}$. The set contains $n+1$ stationary points, all of which are isolated due to Assumption 2. Geometrically, the undesired stationary point $\mathbf{s}_{i}$ (on the boundary of $\mathcal{O}_{i}^{\epsilon}$) and $\mathbf{x}^{\ast}$ are collinear with $\mathbf{c}_{i}$, but located on the opposite sides of $\mathbf{c}_{i}$, as seen in Fig. 13. As per (44), the analytical form of $\mathbf{s}_{i}$ is obtained as follows:
@@
The isolated stationary points $\mathbf{s}_{i}$, $i\in\mathbb{I}$ may prevent us from achieving the objective (motion-to-goal). To check the stability of $\mathbf{s}_{i}$, we define a small neighborhood of $\mathbf{s}_{i}$, i.e., an open ball $\mathcal{B}(\mathbf{s}_{i},\epsilon^{\prime})$ with $0<\epsilon^{\prime}\leq\min\{\epsilon,\epsilon^{\ast}-\epsilon\}$, such that $\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{b}(\mathbf{x})>0$ holds for all $\mathbf{x}\in\mathcal{B}(\mathbf{s}_{i},\epsilon^{\prime})$. When restricted on $\mathcal{B}(\mathbf{s}_{i},\epsilon^{\prime})$, the feasible set of initial configurations is $\mathcal{F}_{i}:=\mathcal{X}_{\epsilon}\cap\mathcal{B}(\mathbf{s}_{i},\epsilon^{\prime})$, and $\mathbf{x}$ evolves according to
$\dot{\mathbf{x}}=-k_{0}(\mathbf{I}_{2}-\phi(\text{d}_{\mathcal{O}}(\mathbf{x}))\mathbf{b}(\mathbf{x})\mathbf{b}^{\top}(\mathbf{x}))(\mathbf{x}-\mathbf{x}^{\ast}).$ (47)
@@
Geometrically speaking, all elements of $\mathcal{M}_{i}$ form a radial line extending outward from the stationary point $\mathbf{s}_{i}$ along $\mathbf{s}_{i}-\mathbf{c}_{i}$ (see the yellow line in Fig. 13), while all elements of $\mathcal{N}_{i}$ form a curve, which is a portion of the boundary $\partial\mathcal{O}_{i}^{\epsilon}$ (see the blue curve in Fig. 13). In the following, three cases are considered: 1) $\mathbf{x}(0)\in\mathcal{B}(\mathbf{s}_{i},\epsilon^{\prime})\cap\mathcal{M}_{i}$; 2) $\mathbf{x}(0)\in\mathcal{N}_{i}\setminus\{\mathbf{s}_{i}\}$; and 3) $\mathbf{x}(0)\in\mathcal{F}_{i}\setminus(\mathcal{M}_{i}\cup\mathcal{N}_{i})$.
Case 1): $\mathbf{x}(0)\in\mathcal{B}(\mathbf{s}_{i},\epsilon^{\prime})\cap\mathcal{M}_{i}$. Note that for all $\mathbf{x}\in\mathcal{M}_{i}$, $\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{b}(\mathbf{x})/\|\bm{\kappa}_{0}(\mathbf{x})\|=1$ holds, whereby (47) reduces to
$\dot{\mathbf{x}}=-k_{0}(1-\phi(\text{d}_{\mathcal{O}}(\mathbf{x})))(\mathbf{x}-\mathbf{x}^{\ast}).$ (50)
From (13) and (48), it is evident that $0\leq\phi(\text{d}_{\mathcal{O}}(\mathbf{x}))\leq 1$ on the set $\mathcal{M}_{i}$ and $\phi(\text{d}_{\mathcal{O}}(\mathbf{x}))=1$ only when $\mathbf{x}=\mathbf{s}_{i}$. Hence, for all $\mathbf{x}\in\mathcal{M}_{i}\setminus\{\mathbf{s}_{i}\}$, the velocity vector $\mathbf{h}(\mathbf{x})$ (i.e., the right-hand side of (50)) points directly towards $\mathbf{s}_{i}$ and becomes zero when $\mathbf{x}=\mathbf{s}_{i}$, indicating that the set $\mathcal{M}_{i}$ is forward invariant. Consider the following LFC: